The perimeter of a rectangle is 50 yds. What are the dimensions that will produce the maximum area of such a rectangle?
Answer by checkley77(12844)

THE MAX. AREA WOULD BE A SQUARE.
50/4=12.5^2=156.25 YDS.^2
PROOF:
12*13=156 YDS^2
12.4*12.6=156.24 YDS^2